# Metric Distortion of Line-up Elections: The Right Person for the Right Job

### v1

@@ Abstract
We provide mechanisms and new metric distortion bounds for line-up elections. In such elections, a set of $n$ voters, $k$ candidates, and $\ell$ positions are all located in a metric space. The goal is to choose a set of candidates and assign them to different positions, so as to minimize the total cost of the voters. The cost of each voter consists of the distances from itself to the chosen candidates (measuring how much the voter likes the chosen candidates, or how similar it is to them), as well as the distances from the candidates to the positions they are assigned to (measuring the fitness of the candidates for their positions). Our mechanisms, however, do not know the exact distances, and instead produce good outcomes while only using a smaller amount of information, resulting in small distortion.
@@ 1 Introduction
Previously, most work on metric distortion has considered selecting a single candidate [Abramowitz et al., 2019; Anagnostides et al., 2022; Anshelevich and Postl, 2017; Anshelevich et al., 2018; Anshelevich and Zhu, 2021; Charikar and Ramakrishnan, 2022; Ebadian et al., 2022; Ghodsi et al., 2019; Gkatzelis et al., 2020; Gross et al., 2017; Kizilkaya and Kempe, 2022; Pierczynski and Skowron, 2019; Anshelevich et al., 2024], or selecting a committees of $k$ candidates [Caragiannis et al., 2022b]. In this work we instead consider line-up elections [Boehmer et al., 2020]. In such elections, there are multiple positions that need to be filled (e.g., different spots on a sports team, different duties in a club or a committee, etc.), and a shared pool of available candidates to fill them. The goal is to choose which candidate will be assigned to each position, with the constraint that a candidate can only hold a single position. Not all candidates are equally fit for every position, as some may be more qualified for some positions over others. Thus unlike in the work mentioned above, where we only cared about the total distance from voters to the chosen candidates, we are now concerned both about choosing candidates who are qualified for the position, and also about what candidate best represents the views of the voters.
These types of considerations also occur in facility location settings. For example, we may have several different building sites, and plan to build different types of facilities at some of them (e.g., a post office at one, a food bank at another, etc.). When choosing which sites to select for building, we want to choose ones which are close to the customers. On the other hand, some sites may be more suitable for particular facilities, since they are closer to the places where they would get their supplies (see Figure 1). In fact, our model is actually more general: positions can also be (not necessarily disjoint) sets in the metric space, instead of single points. If the distance to a set $S$ is defined to be the average distance to the items in the set, then all of our results hold for this case. See Appendix A for details and for further motivating examples.
@@
More formally, we have sets of $n$ voters $\mathcal{V}$, $k$ candidates $\mathcal{C}$, and $\ell$ positions $\mathcal{P}$ located in some metric space with distance function $d\colon\mathcal{V}\cup\mathcal{C}\cup\mathcal{P}\to\mathbb{R}^{+}$: distances between voters and candidates represent how much the voter likes them, and distances between candidates and positions represent how qualified the candidate is for this position. The goal is to form a matching so that every position has a candidate assigned to it, and every candidate is assigned to at most one position. How much a voter $v\in\mathcal{V}$ likes when a candidate $c\in\mathcal{C}$ is assigned to position $p\in\mathcal{P}$ depends on both the similarity between $v$ and $c$, as well as $c$’s fitness for the position $p$, and so is measured by $d(v,c)+d(c,p)$: the smaller this quantity the happier voter $v$ becomes. Thus the total cost for a voter is the sum of this quantity over all positions and candidates assigned to those positions, and the Social Cost for this model is defined as the sum of all voter costs, as in all previous work mentioned above.
@@ 1.1 Our Contributions
Voter Preferences Position Preferences
$k$ 1 3 (3) $k$ 1 3 (3)
$k$ $k$ 3 $k$ $k$ $3-\frac{1}{2^{k-2}}$
[…]
Position Preferences with Voter Preferences
2 1 2 (2)
Candidate and Position Locations Candidate and Position Locations with Voter Preferences
$k$ $1$ 3 (3) 2 1 $\frac{5}{3}$ ($\frac{5}{3}$)
$k$ $k$ 1 $k$ 1 $\frac{25}{9}$
[…]
@@
Table 1: Summary of results showing upper/lower bounds on distortion for lineup elections with different sets of information. The first column is the number of candidates, the second is the number of positions, and the third shows the distortion bounds. Numbers in parentheses are lower bounds.
@@
In Section 3 we consider the classic notion of distortion, where only the ordinal voter preferences are known, i.e., we know the order of which candidates each voter prefers for each position. It is not difficult to show that for ($k$,$1$)-line-up elections, i.e., when there is only a single position, our model is equivalent to usual elections where a single candidate is chosen. Thus existing results for distortion in social choice apply for this case. For the special case of ($k$,$k$)-line-up elections we can also show a distortion of at most 3: this case is simpler than the general one since all candidates must be selected, and we only need to form an appropriate matching without knowing the true distances between the candidates and positions. One of our main results is combining techniques for social choice and matching to give a distortion bound of 7 for general ($k$$\ell$)-line-up elections. The mechanism to achieve this bound is simple: it is an iterative election which chooses a candidate to assign to each position separately by running a single-winner election using the voter preferences. Proving the distortion bound, however, requires a careful charging argument between the costs of various components of the matching between candidates and positions, and the distances from voters to candidates. We first show a simpler distortion bound of 9 to build intuition, and then improve it to a bound of 7 by more carefully bounding the combination of loss from choosing the incorrect matching, and choosing the incorrect set of candidates to match to positions.
@@
We then consider different types of information. In Section 4, we study mechanisms which know ordinal information about the fit of the candidates for each position (which we refer to as “position preferences”). In social choice, it is often reasonable to know the relative fitness of each candidate for each position (e.g., how much experience they have or their qualifications), even if we don’t know the exact fitness levels or the voter preferences. Perhaps surprisingly, this small amount of ordinal information (as compared to knowing the preferences of all the voters) is enough to provide better distortion than if we knew all voter preferences. In fact, for ($k$,$k$)-line-up elections, we can form distortion better than the best possible distortion for standard (non-line-up) elections, and for general ($k$$\ell$)-line-up elections we can improve our bound from 7 to 5. Once again, the mechanism achieving these bounds is simple: it is serial dictatorship by using the position preferences. We are able to achieve these better results by taking existing bounds on the quality of matching and of social choice, and showing tradeoffs between them which combine to form better results than they can separately.
@@
Finally, in Section 5, we consider the case where instead we are given the exact locations of all candidates and positions (and thus we know the distances between them), but we know absolutely nothing about the voter preferences. For example, in Figure 1 we may not know the identities or locations of the stores, but have full information about where products are produced and where the distribution centers are located. This information is enough to form a distortion of 3 for ($k$$\ell$)-line-up elections, even without knowing anything about voters or their preferences. Interestingly, if we also know the ordinal preferences of the voters, then for a single position we can improve distortion to $\frac{5}{3}$ for ($2$,$1$)-line-up elections and to $\frac{25}{9}$ for ($k$,$1$)-line-up elections. Contrast this with distortion for standard single-winner elections, where knowing the exact locations of all the candidates in addition to ordinal voter preferences does not improve the possible distortion as compared to only knowing voter preferences Anshelevich and Zhu [2021].
@@ 1.2 Related Work
Mechanisms for minimizing distortion, i.e., forming mechanisms with access to only ordinal information which perform almost as well as mechanisms with full information, have received a lot of attention. See Anshelevich, Filos-Ratsikas, Shah, and Voudouris [2021] for a survey on this topic. Much of this work specifically focuses on metric distortion, i.e., on forming such mechanisms with the assumption that voters and candidates are located in some unknown metric space (see e.g., Anagnostides et al. [2022]; Anshelevich and Postl [2017]; Anshelevich et al. [2018]; Anshelevich and Zhu [2021]; Caragiannis et al. [2022b]; Charikar and Ramakrishnan [2022]; Cheng et al. [2017]; Feldman et al. [2016]; Ghodsi et al. [2019]; Goel et al. [2017]; Pierczynski and Skowron [2019]; Skowron and Elkind [2017]; Anshelevich et al. [2024]; Kalayci et al. [2024]). Most of this work concerns single-winner elections, with Gkatzelis et al. [2020]; Kizilkaya and Kempe [2022] showing how to achieve a metric distortion of 3 via a deterministic mechanism in single-winner elections. Few papers have looked at distortion of multi-winner elections, with the following main exceptions. Chen et al. [2020] look at selecting a committee equal in size to the total number of candidates minus one. They show a tight bound of 3 on distortion for the case when the cost of a voter equals the distance to the closest member of the committee. Caragiannis et al. [2022b] also consider the problem of picking a committee of $k$ candidates, but such that the cost for each voter is given by the distance to the $q$’th closest candidate. They then try to minimize the sum of the costs of all agents, and show that the distortion may become arbitrarily large when $q\leq\frac{k}{3}$, asymptotically linear in the number of agents when $\frac{k}{3}<q\leq\frac{k}{2}$ and constant when $q>\frac{k}{2}$. Our model differs from that of Caragiannis et al. [2022b] in a few ways. First, we must both choose $k$ candidates to select, as well as which of these is assigned to which position. Second, the cost of a voter is the sum of its costs for all positions, so it is the sum of distances to the selected candidates together with their distances to the assigned positions. Thus, our model is both more difficult (because we must form an assignment of candidates to positions as well), and simpler (since we are looking at the sum of distances instead of the $q$’th closest distance). Note that if there were no positions and we simply wished to choose $k$ candidates $K\subseteq\mathcal{C}$ to minimize $\sum_{v\in\mathcal{V}}\sum_{c\in K}d(v,c)$, then it is easy to achieve a distortion of 3 by using plurality veto [Kizilkaya and Kempe, 2022] to select $k$ candidates one at a time, as in Goel et al. [2018].
@@ 2 Model and Preliminaries
We now formally introduce the ($k$$\ell$)-line-up election and define some terms that will be useful later on in this paper. Let $\mathcal{P}$ be a set of $\ell$ positions, $\mathcal{C}$ be a set of $k$ candidates, and $\mathcal{V}$ be the set of $n$ voters in a metric space with distance function $d$. Throughout this paper we will assume $\lvert\mathcal{P}\rvert\leq\lvert\mathcal{C}\rvert$. The goal in a line-up election is to form an assignment from the set of positions to the set of candidates, so that each candidate is assigned to at most one position, and each position has a candidate assigned to it. In other words, the goal is to form an injective matching $M\colon\mathcal{P}\to\mathcal{C}$. If we assign position $p$ to candidate $c$, then the cost for voter $v$ is given by $d(v,c)+d(c,p)$; this means that the voter cares both about her proximity to the chosen candidate as well as the candidate’s fitness for the position. Thus the total cost of a voter $v$ for a matching $M$ is defined as the following.
@@
In later sections, we also consider mechanisms with ordinal knowledge of how qualified different candidates are for each position. In other words, we know, for each position $p$ and candidates $c_{1},c_{2}$, whether $d(c_{1},p)$ or $d(c_{2},p)$ is smaller, but not the actual distance values. It is often less difficult to sort candidates in order of their qualifications for a position, instead of figuring out the exact numerical level of fitness. We refer to this knowledge as “position preferences” for the candidates.
@@ 3 Knowing Ordinal Voter Preferences
For each ($k$,$1$)-line-up election there exists a standard election with the same candidate set $\mathcal{C}$, such that choosing $c\in\mathcal{C}$ in the line-up election is equal to the cost of choosing $c$ in the standard election.
@@
For any mechanism that achieves a distortion of $\delta$ for the standard election there exists a mechanism that achieves a distortion of $\delta$ for the ($k$,$1$)-line-up election.
@@
By Lemma 3.1, we know that under any metric space $d$ for the ($k$,$1$)-line-up election, there exists a metric space $d^{\prime}$ such that for each candidate $c\in\mathcal{C}$ the corresponding candidate in $d^{\prime}$ has the same cost as $c$ in the line-up election. In addition, the preferences of voters must stay the same since $d^{\prime}(v,c)=d(v,c)+d(c,p)$ for position $p$ and voter $v$. Then, to solve the original ($k$,$1$)-line-up election, we can instead run a standard election with candidates $\mathcal{C}$ and the same voter preferences. If we choose candidate $c$ for the standard election we obtain the same cost for the line-up election by choosing $c$ for the position. Therefore, if we have a mechanism that achieves a distortion of $\delta$ of the standard election, by using the reduction above, for the resulting solution $c$ and the optimal solution $c^{*}$, we must have $\sum_{v\in\mathcal{V}}\left[d(v,c)+d(c,p)\right]\leq\delta\sum_{v\in\mathcal{V}}\left[d(v,c^{*})+d(c^{*},p)\right]$. ∎
@@
There exists a mechanism for the the ($k$,$1$)-line-up election that achieves a distortion of $3$. Moreover, no other deterministic mechanism can achieve a worst-case distortion bound less than 3, even on a line with 2 candidates.
@@
From Gkatzelis et al. [2020]; Kizilkaya and Kempe [2022] we know that there exists a deterministic mechanism to solve the standard election which achieves a distortion of $3$. Then, by Lemma 3.2, we can conclude that there exists a mechanism for the ($k$,$1$)-line-up election that achieves a distortion of $3$.
@@
Therefore, by Theorem 3.3 we know that we can solve ($k$,$1$)-line-up election and achieve the optimal distortion of $3$ using standard techniques from metric distortion. However, in Lemma 3.2 we also consider distortions other than 3 for general mechanisms since mechanisms such as plurality veto, which achieves optimal distortion, may not have desirable properties such as strategy-proofness. In addition, the mechanism we chose may also be given additional information beyond what we have assumed so far in this section, which would lead to distortions other than three, for example when information about the decisiveness of voters Anshelevich and Postl [2017] or when information about each voter’s passion for each candidate pair is known Abramowitz et al. [2019]. Now that we have discussed results for the case where there is only one position, we will then look at cases where we have more than one position. Before we look at the general case where we have arbitrary $\ell$ positions and $k$ candidates, we first consider the special case where $\ell=k$.
@@
For any ($k$,$k$)-line-up election all matchings are at most a factor of 3 away from optimal.
@@
After having looked at special cases of $(k,1)$ and ($k$,$k$)-line-up elections, we will now look at a mechanism for the more general ($k$$\ell$)-line-up election problem. We introduce iterative election, a natural mechanism for the ($k$$\ell$)-line-up election problem. For a position $p$ reduce the problem to a ($k$,$1$)-line-up election problem using the transformation given in Lemma 3.1, and then use a mechanism for the standard election problem. If $c$ is selected using the standard election mechanism we assign $M_{1}(p)=c$. We then remove the selected candidate $c$, and repeat the procedure for each remaining position, removing the selected candidate each time. Importantly, this mechanism is quite general, as we do not assume anything about the mechanism or the specific information set $I$. One could achieve optimal metric distortion for each position by using plurality veto Kizilkaya and Kempe [2022] for instance, or use a different mechanism that has attractive properties like strategyproofness. Lastly, iterative election is also an online algorithm, meaning that positions can arrive overtime, and as long as we have the appropriate information we can immediately assign each position at the time of arrival. We will now analyze the case where we utilize a standard election mechanism with distortion $\delta_{itr}$ and show an upper bound for the iterative election procedure on the ($k$$\ell$)-line-up election problem.
@@
We also note that $M_{2}(M_{1}\raisebox{4.95134pt}{$\scriptscriptstyle-\!1$})$ is a bijection from $S_{1}$ to $S_{3}$. Thus we can sum over $c^{\prime}=M_{2}(M_{1}\raisebox{4.95134pt}{$\scriptscriptstyle-\!1$}(c))$ in $S_{3}$ instead of over $S_{1}$ and obtain

### v2

@@ Abstract
We provide mechanisms and new metric distortion bounds for line-up elections. In such elections, a set of $n$ voters, $m$ candidates, and $\ell$ positions are all located in a metric space. The goal is to choose a set of candidates and assign them to different positions, so as to minimize the total cost of the voters. The cost of each voter consists of the distances from itself to the chosen candidates (measuring how much the voter likes the chosen candidates, or how similar it is to them), as well as the distances from the candidates to the positions they are assigned to (measuring the fitness of the candidates for their positions). Our mechanisms, however, do not know the exact distances, and instead produce good outcomes while only using a smaller amount of information, resulting in small distortion.
@@ 1 Introduction
Previously, most work on metric distortion has considered selecting a single candidate [Abramowitz et al., 2019; Anagnostides et al., 2022; Anshelevich and Postl, 2017; Anshelevich et al., 2018; Anshelevich and Zhu, 2021; Charikar and Ramakrishnan, 2022; Ebadian et al., 2022; Ghodsi et al., 2019; Gkatzelis et al., 2020; Gross et al., 2017; Kizilkaya and Kempe, 2022; Pierczynski and Skowron, 2019; Anshelevich et al., 2024], or selecting a committees of $m$ candidates [Caragiannis et al., 2022b]. In this work we instead consider line-up elections [Boehmer et al., 2020]. In such elections, there are multiple positions that need to be filled (e.g., different spots on a sports team, different duties in a club or a committee, etc.), and a shared pool of available candidates to fill them. The goal is to choose which candidate will be assigned to each position, with the constraint that a candidate can only hold a single position. Not all candidates are equally fit for every position, as some may be more qualified for some positions over others. Thus, unlike in the work mentioned above, where we only cared about the total distance from voters to the chosen candidates, we are now concerned both about choosing candidates who are qualified for the position and also about what candidate best represents the views of the voters.
These types of considerations also occur in facility location settings. For example, we may have several different building sites and plan to build different types of facilities at some of them (e.g., a post office at one, a food bank at another, etc.). When choosing which sites to select for building, we want to choose ones which are close to the customers. On the other hand, some sites may be more suitable for particular facilities, since they are closer to the places where they would get their supplies (see Figure 1). In fact, our model is actually more general: positions can also be (not necessarily disjoint) sets in the metric space, instead of single points. If the distance to a set $S$ is defined to be the average distance to the items in the set, then all of our results hold for this case. See Appendix A for details and for further motivating examples.
@@
More formally, we have sets of $n$ voters $\mathcal{V}$, $m$ candidates $\mathcal{C}$, and $\ell$ positions $\mathcal{P}$ located in some metric space with distance function $d\colon\mathcal{V}\cup\mathcal{C}\cup\mathcal{P}\to\mathbb{R}^{+}$: distances between voters and candidates represent how much the voter likes them, and distances between candidates and positions represent how qualified the candidate is for this position. The goal is to form a matching so that every position has a candidate assigned to it, and every candidate is assigned to at most one position. How much a voter $v\in\mathcal{V}$ likes when a candidate $c\in\mathcal{C}$ is assigned to position $p\in\mathcal{P}$ depends on both the similarity between $v$ and $c$, as well as $c$”s fitness for position $p$, and is therefore measured by $d(v,c)+d(c,p)$: the smaller this quantity, the happier the voter $v$ becomes. Thus, the total cost for a voter is the sum of this quantity over all positions and candidates assigned to those positions, and the Social Cost for this model is defined as the sum of all voter costs, as in all previous work mentioned above.
@@
Importantly, a limitation of our model is that we require the sets $\mathcal{V},\mathcal{P}$, and $\mathcal{C}$ to all lie in a common metric space. This is a crucial requirement: if we allow for multiple arbitrary metric spaces then it is easy to show that in the worst case our model cannot give bounds better than the standard matching model, which has exponential distortion [Caragiannis et al., 2022a]. That said, there are numerous settings where this assumption holds, such as in distribution networks where all locations $\mathcal{P}$ creating the goods, locations $\mathcal{V}$ consuming goods, and intermediate distribution centers $\mathcal{C}$ all lie in a common space. For more discussion and further examples see the full version of this paper.
@@ 1.1 Our Contributions
$(m,1)$ $(m,m)$ $(m,\ell)$ (2,1)
VP 3 (3) 3 7 (3) 3 (3)
PP 3 (3) $3-\frac{1}{2^{m-2}}$ 5 (3) 3 (3)
LOC 3 (3) 1 3 (3) 3 (3)
VP+PP - - - 2 (2)
VP+LOC $\frac{25}{9}$ 53($\frac{5}{3}(\frac{5}{3})$
@@
Table 1: Summary of results showing upper/lower bounds on distortion for lineup elections with different types of available information. VP stands for mechanisms that operate knowing voter preferences, PP for position preferences, and LOC for candidate and position locations. Numbers in parentheses are lower bounds.
@@
In Section 3 we consider the classic notion of distortion, where only the ordinal voter preferences are known, i.e., we know the order of which candidates each voter prefers for each position. It is not difficult to show that for ($m$,$1$)-line-up elections, i.e., when there is only a single position, our model is equivalent to usual elections where a single candidate is chosen. Thus existing results for distortion in social choice apply for this case. For the special case of ($m$,$m$)-line-up elections we can also show a distortion of at most 3: this case is simpler than the general one since all candidates must be selected, and we only need to form an appropriate matching without knowing the true distances between the candidates and positions. One of our main results is combining techniques for social choice and matching to give a distortion bound of 7 for general ($m$$\ell$)-line-up elections. The mechanism to achieve this bound is simple: it is an iterative election which chooses a candidate to assign to each position separately by running a single-winner election using the voter preferences. Proving the distortion bound, however, requires a careful charging argument between the costs of various components of the matching between candidates and positions, and the distances from voters to candidates. We first show a simpler distortion bound of 9 to build intuition, and then improve it to a bound of 7 by more carefully bounding the combination of loss from choosing the incorrect matching, and choosing the incorrect set of candidates to match to positions.
@@
We then consider different types of information. In Section 4, we study mechanisms which know ordinal information about the fit of the candidates for each position (which we refer to as “position preferences”). In social choice, it is often reasonable to know the relative fitness of each candidate for each position (e.g., how much experience they have or their qualifications), even if we don’t know the exact fitness levels or the voter preferences. Perhaps surprisingly, this small amount of ordinal information (as compared to knowing the preferences of all the voters) is enough to provide better distortion than if we knew all voter preferences. In fact, for ($m$,$m$)-line-up elections, we can form distortion better than the best possible distortion for standard (non-line-up) elections, and for general ($m$$\ell$)-line-up elections we can improve our bound from 7 to 5. Once again, the mechanism achieving these bounds is simple: it is serial dictatorship by using the position preferences. We are able to achieve these better results by taking existing bounds on the quality of matching and of social choice, and showing trade-offs between them which combine to form better results than they can separately.
@@
Finally, in Section 5, we consider the case where instead we are given the exact locations of all candidates and positions (and thus we know the distances between them), but we know absolutely nothing about the voter preferences. For example, in Figure 1 we may not know the identities or locations of the stores, but have full information about where products are produced and where the distribution centers are located. This information is enough to form a distortion of 3 for ($m$$\ell$)-line-up elections, even without knowing anything about voters or their preferences. Interestingly, if we also know the ordinal preferences of the voters, then for a single position we can improve distortion to $\frac{5}{3}$ for ($2$,$1$)-line-up elections and to $\frac{25}{9}$ for ($m$,$1$)-line-up elections. Contrast this with distortion for standard single-winner elections, where knowing the exact locations of all the candidates in addition to ordinal voter preferences does not improve the possible distortion as compared to only knowing voter preferences Anshelevich and Zhu [2021].
@@ 1.2 Related Work
Mechanisms for minimizing distortion, i.e., forming mechanisms with access to only ordinal information which perform almost as well as mechanisms with full information, have received a lot of attention. See Anshelevich, Filos-Ratsikas, Shah, and Voudouris [2021] for a survey on this topic. Much of this work specifically focuses on metric distortion, i.e., on forming such mechanisms with the assumption that voters and candidates are located in some unknown metric space (see e.g., Anagnostides et al. [2022]; Anshelevich and Postl [2017]; Anshelevich et al. [2018]; Anshelevich and Zhu [2021]; Caragiannis et al. [2022b]; Charikar and Ramakrishnan [2022]; Cheng et al. [2017]; Feldman et al. [2016]; Ghodsi et al. [2019]; Goel et al. [2017]; Pierczynski and Skowron [2019]; Skowron and Elkind [2017]; Anshelevich et al. [2024]; Kalayci et al. [2024]). Most of this work concerns single-winner elections, with Gkatzelis et al. [2020]; Kizilkaya and Kempe [2022] showing how to achieve a metric distortion of 3 via a deterministic mechanism in single-winner elections. Few papers have looked at distortion of multi-winner elections, with the following main exceptions. Chen et al. [2020] look at selecting a committee equal in size to the total number of candidates minus one. They show a tight bound of 3 on distortion for the case when the cost of a voter equals the distance to the closest member of the committee. Caragiannis et al. [2022b] also consider the problem of picking a committee of $k$ candidates, but such that the cost for each voter is given by the distance to the $q$’th closest candidate. They then try to minimize the sum of the costs of all agents, and show that the distortion may become arbitrarily large when $q\leq\frac{k}{3}$, asymptotically linear in the number of agents when $\frac{k}{3}<q\leq\frac{k}{2}$ and constant when $q>\frac{k}{2}$. Our model differs from that of Caragiannis et al. [2022b] in a few ways. First, we must both choose $m$ candidates to select, as well as which of these is assigned to which position. Second, the cost of a voter is the sum of its costs for all positions, so it is the sum of distances to the selected candidates together with their distances to the assigned positions. Thus, our model is both more difficult (because we must form an assignment of candidates to positions as well), and simpler (since we are looking at the sum of distances instead of the $q$’th closest distance). Note that if there were no positions and we simply wished to choose $k$ candidates $K\subseteq\mathcal{C}$ to minimize $\sum_{v\in\mathcal{V}}\sum_{c\in K}d(v,c)$, then it is easy to achieve a distortion of 3 by using plurality veto [Kizilkaya and Kempe, 2022] to select $k$ candidates one at a time, as in Goel et al. [2018].
@@ 2 Model and Preliminaries
We now formally introduce the ($m$$\ell$)-line-up election and define some terms that will be useful later on in this paper. Let $\mathcal{P}$ be a set of $\ell$ positions, $\mathcal{C}$ be a set of $m$ candidates, and $\mathcal{V}$ be the set of $n$ voters in a metric space with distance function $d$. Throughout this paper we will assume $\lvert\mathcal{P}\rvert\leq\lvert\mathcal{C}\rvert$. The goal in a line-up election is to form an assignment from the set of positions to the set of candidates, so that each candidate is assigned to at most one position, and each position has a candidate assigned to it. In other words, the goal is to form an injective matching $M\colon\mathcal{P}\to\mathcal{C}$. If we assign position $p$ to candidate $c$, then the cost for voter $v$ is given by $d(v,c)+d(c,p)$; this means that the voter cares both about her proximity to the chosen candidate as well as the candidate’s fitness for the position. Thus the total cost of a voter $v$ for a matching $M$ is defined as the following.
@@
In later sections, we also consider mechanisms with ordinal knowledge of how qualified different candidates are for each position. In other words, we know, for each position $p$ and candidates $c_{1},c_{2}$, whether $d(c_{1},p)$ or $d(c_{2},p)$ is smaller, but not the actual distance values. It is often less difficult to sort candidates in order of their qualifications for a position, instead of figuring out the exact numerical level of fitness. We refer to this knowledge as ”position preferences” for the candidates.
@@ 3 Knowing Ordinal Voter Preferences
For each ($m$,$1$)-line-up election there exists a standard election with the same candidate set $\mathcal{C}$, such that choosing $c\in\mathcal{C}$ in the line-up election is equal to the cost of choosing $c$ in the standard election.
@@
For any mechanism that achieves a distortion of $\delta$ for the standard election there exists a mechanism that achieves a distortion of $\delta$ for the ($m$,$1$)-line-up election.
@@
By Lemma 3.1, we know that under any metric space $d$ for the ($m$,$1$)-line-up election, there exists a metric space $d^{\prime}$ such that for each candidate $c\in\mathcal{C}$ the corresponding candidate in $d^{\prime}$ has the same cost as $c$ in the line-up election. In addition, the preferences of voters must stay the same since $d^{\prime}(v,c)=d(v,c)+d(c,p)$ for position $p$ and voter $v$. Then, to solve the original ($m$,$1$)-line-up election, we can instead run a standard election with candidates $\mathcal{C}$ and the same voter preferences. If we choose candidate $c$ for the standard election we obtain the same cost for the line-up election by choosing $c$ for the position. Therefore, if we have a mechanism that achieves a distortion of $\delta$ of the standard election, by using the reduction above, for the resulting solution $c$ and the optimal solution $c^{*}$, we must have $\sum_{v\in\mathcal{V}}\left[d(v,c)+d(c,p)\right]\leq\delta\sum_{v\in\mathcal{V}}\left[d(v,c^{*})+d(c^{*},p)\right]$. ∎
@@
There exists a mechanism for the the ($m$,$1$)-line-up election that achieves a distortion of $3$. Moreover, no other deterministic mechanism can achieve a worst-case distortion bound less than 3, even on a line with 2 candidates.
@@
From Gkatzelis et al. [2020]; Kizilkaya and Kempe [2022] we know that there exists a deterministic mechanism to solve the standard election which achieves a distortion of $3$. Then, by Lemma 3.2, we can conclude that there exists a mechanism for the ($m$,$1$)-line-up election that achieves a distortion of $3$.
@@
Therefore, by Theorem 3.3 we know that we can solve ($m$,$1$)-line-up election and achieve the optimal distortion of $3$ using standard techniques from metric distortion. However, in Lemma 3.2 we also consider distortions other than 3 for general mechanisms since mechanisms such as plurality veto, which achieves optimal distortion, may not have desirable properties such as strategy-proofness. In addition, the mechanism we chose may also be given additional information beyond what we have assumed so far in this section, which would lead to distortions other than three, for example when information about the decisiveness of voters Anshelevich and Postl [2017] or when information about each voter’s passion for each candidate pair is known Abramowitz et al. [2019]. Now that we have discussed results for the case where there is only one position, we will then look at cases where we have more than one position. Before we look at the general case where we have arbitrary $\ell$ positions and $m$ candidates, we first consider the special case where $\ell=m$.
@@
For any ($m$,$m$)-line-up election all matchings are at most a factor of 3 away from optimal.
@@
After having looked at special cases of $(m,1)$ and ($m$,$m$)-line-up elections, we will now look at a mechanism for the more general ($m$$\ell$)-line-up election problem. We introduce iterative election, a natural mechanism for the ($m$$\ell$)-line-up election problem. For a position $p$ reduce the problem to a ($m$,$1$)-line-up election problem using the transformation given in Lemma 3.1, and then use a mechanism for the standard election problem. If $c$ is selected using the standard election mechanism we assign $M_{1}(p)=c$. We then remove the selected candidate $c$, and repeat the procedure for each remaining position, removing the selected candidate each time. Importantly, this mechanism is quite general, as we do not assume anything about the mechanism or the specific information set $I$. One could achieve optimal metric distortion for each position by using plurality veto Kizilkaya and Kempe [2022] for instance, or use a different mechanism that has attractive properties like strategyproofness. Lastly, iterative election is also an online algorithm, meaning that positions can arrive overtime, and as long as we have the appropriate information, we can immediately assign each position at the time of arrival. We will now analyze the case where we utilize a standard election mechanism with distortion $\delta_{itr}$ and show an upper bound for the iterative election procedure on the ($m$$\ell$)-line-up election problem.
@@
We also note that $M_{2}(M_{1}\raisebox{4.95134pt}{$\scriptscriptstyle-\!1$})$ is a bijection from $S_{1}$ to $S_{3}$. Thus, we can sum over $c^{\prime}=M_{2}(M_{1}\raisebox{4.95134pt}{$\scriptscriptstyle-\!1$}(c))$ in $S_{3}$ instead of over $S_{1}$ and obtain